The Kim Company is a furniture manufacturer with two departments: molding and finishing. The company uses the weighted-average method of process costing. In August, the following data were recorded for the finishing department:

Units of beginning work in process inventory 12,500
Units transfered in during August 87,500
Units completed 62,500
Transferred-in costs in beginning WIP $103,625
Cost of units transferred in during current period $809,375

There were no spoiled or lost units. The firm had completed 24% of the work for the units in ending WIP.

Required: Compute the transferred-in costs contained in the cost of goods manufactured in August.
Answer #1

Correct answer-------------$ 798,077 or $ 798,076.92

Working

Reconciliation of Units
A Beginning WIP                   12,500
B Introduced                   87,500
C=A+B TOTAL                 100,000
D Transferred out                   62,500
E=C-D Ending WIP                   37,500

.

Statement of Equivalent Units(Weighted average)
Transferred in
Units Complete % Equivalent units
Transferred out                                62,500 100%                    62,500
Ending WIP                                37,500 24%                       9,000
Total                             100,000 Total                    71,500

.

Cost per Equivalent Units (Weighted average)
COST Transferred in
Beginning WIP Inventory Cost (A) $ 103,625
Cost incurred during period (B) $ 809,375
Total Cost to be accounted for (C=A+B) $ 913,000
Total Equivalent Units(D)                   71,500
Cost per Equivalent Units (E=C/D) $ 12.77

.

.

Statement of cost (Weighted average)
Cost Equivalent Cost/unit Transferred
Units Cost Allocated
Transfferd $ 12.77                       62,500 $        798,076.92
TOTAL $ 913,000 TOTAL $ 798,077